# The large-$N$ limit of the topological susceptibility of $\mathrm{SU}(N)$ Yang–Mills theories via Parallel Tempering on Boundary Conditions

### v1

@@ Abstract
I present a large-$N$ determination of the topological susceptibility $\chi$ of SU(N)(N) Yang–Mills theories using non-perturbative numerical Monte Carlo simulations of the lattice-discretized theory for $3\leq N\leq 6$, and adopting the Parallel Tempering on Boundary Conditions (PTBC) algorithm to bypass topological freezing for $N>3$. Thanks to this algorithm I am able to explore a uniform range of lattice spacing across all values of $N$, and to precisely determine $\chi$ for finer lattice spacings compared to previous studies with periodic or open boundary conditions. By taking the continuum limit at fixed smoothing radius in physical units, I am also able to show the independence of the continuum limit of $\chi$ from this choice. I conclude providing a comprehensive comparison of my new PTBC results with previous determinations of the topological susceptibility in the literature, both at finite $N$ and in the large-$N$ limit.
@@ 1 Introduction
On general grounds, lattice simulations have proven to be a reliable tool to study large-$N$ gauge theories, and have been extensively employed to this end DelDebbio:2001sj ; Lucini:2001ej ; Lucini:2004my ; DelDebbio:2006yuf ; Vicari:2008jw ; Allton:2008ty ; Lucini:2010nv ; Lucini:2012gg ; Bali:2013kia ; Bonati:2016tvi ; Ce:2016awn ; Bennett:2020hqd ; Athenodorou:2021qvs ; Bennett:2022gdz ; Athenodorou:2024loq ; Sharifian:2025fyl ; DeGrand:2016pur ; Hernandez:2019qed ; DeGrand:2020utq ; Hernandez:2020tbc ; DeGrand:2021zjw ; Baeza-Ballesteros:2022azb ; DeGrand:2023hzz ; DeGrand:2024lvp ; DeGrand:2024frm ; Baeza-Ballesteros:2025iee ; Butti:2025rnu ; PhysRevLett.48.1063 ; BHANOT198247 ; Gross:1982at ; GONZALEZARROYO1983174 ; PhysRevD.27.2397 ; Aldazabal:1983ec ; Kiskis:2002gr ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch ; Gonzalez-Arroyo:2010omx ; Neuberger:2020wpx ; Gonzalez-Arroyo:1983cyv ; Das:1984jh ; Das:1985xc ; Narayanan:2004cp ; Gonzalez-Arroyo:2005dgf ; Kiskis:2009rf ; Hietanen:2009ex ; Hietanen:2010fx ; Bringoltz:2011by ; Hietanen:2012ma ; Gonzalez-Arroyo:2012euf ; Gonzalez-Arroyo:2013bta ; Lohmayer:2013spa ; Gonzalez-Arroyo:2014dua ; GarciaPerez:2014azn ; GarciaPerez:2015rda ; Perez:2015ssa ; Gonzalez-Arroyo:2015bya ; Perez:2017jyq ; GarciaPerez:2020gnf ; Perez:2020vbn ; Butti:2023hfp ; Bonanno:2023ypf ; Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr ; Hamada:2025whg ; Bonanno:2025hzr . However, the lattice investigation of gauge topology in the large-$N$ limit is known to be a challenging task due to a serious algorithmic limitation: the infamous topological freezing problem. When adopting standard local updating algorithms to sample the lattice Yang–Mills path integral, the autocorrelation time of $Q$, i.e., the number of MC steps necessary to draw two decorrelated occurrences of the topological charge, rapidly diverges as function of the inverse lattice spacing $a$ as the continuum limit $a\to 0$ is approached Alles:1996vn ; DelDebbio:2002xa ; DelDebbio:2004xh ; Schaefer:2010hu . This is understood in terms of the Monte Carlo dynamics of topological modes: achieving a jump of the topological charge of the lattice gauge configuration via a local field deformation becomes increasingly more difficult as the lattice spacing gets finer due to the development of potential barriers among pseudo-topological sectors, which eventually diverge in the continuum limit to restore a proper notion of integer-valued winding number Luscher:1981zq . The rapid growth of the autocorrelation time of the topological charge thus induces a severe loss of ergodicity of the Monte Carlo Markov chain, which tends to remain frozen in a fixed topological sector for a very long time, preventing a correct sampling of the topological charge distribution. Moreover, increasing the rank of the gauge group $N$ at fixed lattice spacings further worsens the severity of freezing: in practice, at large $N$ topological freezing kicks in already at coarse lattice spacing, making it very challenging to study the continuum limit of topological quantities in this regime.
@@
While this algorithm has allowed the determination of several quantities whose calculation was either unsatisfactorily or at all attempted previously with other algorithms Hasenbusch:2017unr ; Berni:2019bch ; Bonanno:2022hmz ; Bonanno:2020hht ; Bonanno:2022yjr ; Bonanno:2023hhp ; Bonanno:2024nba ; Bonanno:2024ggk , so far only a few determinations of $\chi$ using the PTBC algorithm were provided. Some for $N=4,6$ can be found in Ref. Bonanno:2020hht , but no dedicated study of the susceptibility was performed in that paper. The goal of the present investigation is to fill this gap. I will present a dedicated high-statistics lattice study of the topological susceptibility at large $N$ using the PTBC algorithm to effectively mitigate topological freezing. Thanks to PTBC, I will simulate Yang–Mills theories for $N=4,5,6$ across the same range of lattice spacings avoiding topological freezing. The probed parameter space will include finer lattice spacings compared to previous investigations with PBC Bonati:2015sqt ; Bonati:2016tvi ; Athenodorou:2020ani ; Athenodorou:2021qvs and OBC Ce:2015qha ; Ce:2016awn , and will be similar in extension to the one used in typical $N=3$ studies. I will then combine these new large-$N$ results with the the high-statistics ones I obtained for $N=3$ in Bonanno:2023ple to study the large-$N$ limit of $\chi$. Finally, I will conclude with a critical comparison of the determinations of $\chi(N)$ presented in this study with earlier ones obtained with different methods and algorithms.
@@ 2.1 Lattice action and PTBC algorithm
The calculation of $\Delta\mathcal{S}(r)$ has a negligible cost being non-vanishing only around $D$. After swaps, updating sweeps of the full lattice are alternated with some hierarchical updates concentrated around the defect $D$, which is the region where most of the new topological excitations are created/annihilated. These are chosen to have a negligible total numerical cost with respect to the full sweeps (more technical details are found in Hasenbusch:2017unr ; Bonanno:2020hht ). Each hierarchical update is followed by swaps proposals too. Finally, after each series of swap proposals, a random translation of one lattice spacing in a random direction is performed on the periodic lattice, which is translation invariant. Indeed, since the region $D$ is kept fixed, performing random translations of the PBC replica effectively moves the position of the defect, so that topological excitations are created/annihilated around the lattice.
@@ 2.2 Determination of the topological susceptibility
$Q_{\scriptscriptstyle{\mathrm{clov}}}=\frac{1}{32\pi^{2}}\sum_{x}\sum_{\mu\nu\rho\sigma}\varepsilon_{\mu\nu\rho\sigma}\mathrm{Tr}\left[C_{\mu\nu}(x)C_{\rho\sigma}(x)\right]$ (12)
where
@@
It is well-known that gluonic discretizations are affected by multiplicative renormalizations Campostrini:1988cy ; Vicari:2008jw that makes them non-integer on the lattice:
@@
$Q=\int\mathrm{d}^{4}x\,q(x),\qquad q(x)=\frac{1}{16\pi^{2}}\varepsilon_{\mu\nu\rho\sigma}\mathrm{Tr}\left[G_{\mu\nu}(x)G_{\rho\sigma}(x)\right],$ (16)
@@
i.e., one expects that the value of $R_{\scriptscriptstyle{\mathrm{s}}}$ should only influences the slope of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ as a function of $a^{2}$, but not its continuum limit $\chi$. In the following, I will verify this behavior by taking the continuum limit of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ at fixed value of $R_{\scriptscriptstyle{\mathrm{s}}}$ in physical units for several values of $R_{\scriptscriptstyle{\mathrm{s}}}$.
@@ 3 Numerical results
This section is devoted to the presentation of the new PTBC results for the topological susceptibility. I will start by discussing the explored parameter space and the details of the PTBC algorithm, the obtained autocorrelation times and the adopted scale setting. Then, I will present the continuum and large-$N$ extrapolation of the obtained lattice results for $\chi$. Finally, I will compare my new determinations with previous ones in the literature.
@@ 3.1 Simulation parameters, autocorrelation times and scale setting
Indeed, rescaled $\tau_{0}(Q^{2})$ can be fitted to a constant with a reduced chi-squared of $\sim 1.25$ with 10 degrees of freedom, yielding $C^{\prime}=0.080(5)$. This seems to suggest the following combined scaling towards the continuum and large-$N$ limit of the autocorrelation time of $Q^{2}$ in the periodic system:
@@ 3.2 Determination of $\chi$ at large NN
The goal of this section is to discuss the continuum and large-$N$ limit of $\chi$, and to verify the independence of $\chi$ on the choice of the smoothing radius in the continuum limit.
@@
Interestingly, lattice artifacts seems to be practically independent of $N$ for all explored values of $R_{\scriptscriptstyle{\mathrm{s}}}$—i.e., $C(R_{\scriptscriptstyle{\mathrm{s}}},N)\simeq C(R_{\scriptscriptstyle{\mathrm{s}}})$—as revealed by the plots in Fig. 5, where I compare the continuum limits of $\chi_{\scriptscriptstyle{\mathrm{L}}}$ as a function of $N$ at fixed smoothing radius. This fact was already observed in Ce:2016awn . Moreover, lattice artifacts become very small above $R_{\scriptscriptstyle{\mathrm{s}}}\sqrt{\sigma}\gtrsim 1.2$, as it can be seen from Figs. 4 and 5. For example, for the value $R_{\scriptscriptstyle{\mathrm{s}}}\sqrt{\sigma}\simeq 1.223$ reported in those figures, the coarsest lattice spacing determinations of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ differ from their corresponding continuum ones by $\sim 2\%$ at most for all values of $N$. Since the lattice artifacts affecting $\chi_{{\scriptscriptstyle{\mathrm{L}}}}(R_{\scriptscriptstyle{\mathrm{s}}})$ have 3 different sources (the discretized lattice action used in the Monte Carlo, the discretized lattice action used in the smoothing algorithm, and the discretized observable used to compute the topological charge), the flattening of $C(R_{\scriptscriptstyle{\mathrm{s}}})$ for large enough smoothing radii could be due to an accidental cancellation among them. [^1] See Ref. Ramos:2015baa for a similar discussion about the lattice artifacts affecting the energy density computed after the gradient flow within the framework of Symanzik improvement.
@@
NN $\beta$ $a\sqrt{\sigma}$ $a^{4}\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ $\chi_{{\scriptscriptstyle{\mathrm{L}}}}/\sigma^{2}$
@@ 3.4 Comparison with previous determinations of $\chi$ in the literature
0.02651(330) ETMC, Cichy et al. 15 Cichy:2015jra
Wilson (Iwasaki gauge action)
@@
Table 7: Summary of continuum determination of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(3)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 11 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action unless stated otherwise, and all assumed periodic boundaries.
@@
Table 8: Summary of continuum determination of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(4)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 12 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 9: Summary of continuum determination of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(5)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 13 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 10: Summary of continuum determination of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(6)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 14 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 11: Summary of continuum determination of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(\infty)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 15 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
The case $N=3$ has the larger number of entries, and is the only case where I could retrieve determinations obtained both via gluonic definitions Durr:2006ky ; Bonati:2015sqt ; Ce:2015qha ; Athenodorou:2020ani ; Durr:2025qtq —computed employing several smoothing algorithms–and fermionic ones DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg —with calculations performed adopting both chiral and non-chiral quark discretizations. Due to the much larger computational burden, fermionic definitions are systematically less precise compared to gluonic ones. Modulo this caveat, there is overall a good agreement between fermionic and gluonic results, and among all results as a whole. A naive constant fit to the full data set yields a reduced chi-squared of about 1.5 with 9 degrees of freedom, corresponding to a $\sim 15\%$ $p$-value, indeed confirming the observed agreement.

### v2

@@ Abstract
I present a large-$N$ determination of the topological susceptibility $\chi$ of SU(N)(N) Yang–Mills theories using non-perturbative numerical Monte Carlo simulations of the lattice-discretized theory for $3\leq N\leq 6$, and adopting the Parallel Tempering on Boundary Conditions (PTBC) algorithm to bypass topological freezing for $N>3$. Thanks to this algorithm I am able to explore a uniform range of lattice spacings across all values of $N$, and to precisely determine $\chi$ for finer lattice spacings compared to previous studies with periodic or open boundary conditions. By taking the continuum limit at fixed smoothing radius in physical units, I am also able to show the independence of the continuum limit of $\chi$ from this choice. I conclude providing a comprehensive comparison of my new PTBC results with previous determinations of the topological susceptibility in the literature, both at finite $N$ and in the large-$N$ limit.
@@ 1 Introduction
On general grounds, lattice simulations have proven to be a reliable tool to study large-$N$ gauge theories, and have been extensively employed to this end DelDebbio:2001sj ; Lucini:2001ej ; Lucini:2004my ; DelDebbio:2006yuf ; Vicari:2008jw ; Allton:2008ty ; Lucini:2010nv ; Lucini:2012gg ; Bali:2013kia ; Bonati:2016tvi ; Ce:2016awn ; Bennett:2020hqd ; Athenodorou:2021qvs ; Bennett:2022gdz ; Athenodorou:2024loq ; Sharifian:2025fyl ; DeGrand:2016pur ; Hernandez:2019qed ; DeGrand:2020utq ; Hernandez:2020tbc ; DeGrand:2021zjw ; Baeza-Ballesteros:2022azb ; DeGrand:2023hzz ; DeGrand:2024lvp ; DeGrand:2024frm ; Baeza-Ballesteros:2025iee ; Butti:2025rnu ; PhysRevLett.48.1063 ; BHANOT198247 ; Gross:1982at ; GONZALEZARROYO1983174 ; PhysRevD.27.2397 ; Aldazabal:1983ec ; Kiskis:2002gr ; Narayanan:2003fc ; Kovtun:2007py ; Unsal:2008ch ; Gonzalez-Arroyo:2010omx ; Neuberger:2020wpx ; Gonzalez-Arroyo:1983cyv ; Das:1984jh ; Das:1985xc ; Narayanan:2004cp ; Gonzalez-Arroyo:2005dgf ; Kiskis:2009rf ; Hietanen:2009ex ; Hietanen:2010fx ; Bringoltz:2011by ; Hietanen:2012ma ; Gonzalez-Arroyo:2012euf ; Gonzalez-Arroyo:2013bta ; Lohmayer:2013spa ; Gonzalez-Arroyo:2014dua ; GarciaPerez:2014azn ; GarciaPerez:2015rda ; Perez:2015ssa ; Gonzalez-Arroyo:2015bya ; Perez:2017jyq ; GarciaPerez:2020gnf ; Perez:2020vbn ; Butti:2023hfp ; Bonanno:2023ypf ; Butti:2022sgy ; Bonanno:2024bqg ; Bonanno:2024onr ; Hamada:2025whg ; Bonanno:2025hzr . However, the lattice investigation of gauge topology in the large-$N$ limit is known to be a challenging task due to a serious algorithmic limitation: the infamous topological freezing problem. When adopting standard local updating algorithms to sample the lattice Yang–Mills path integral, the autocorrelation time of $Q$, i.e., the number of MC steps necessary to draw two decorrelated occurrences of the topological charge, rapidly diverges as function of the inverse lattice spacing $a$ as the continuum limit $a\to 0$ is approached Alles:1996vn ; DelDebbio:2002xa ; DelDebbio:2004xh ; Schaefer:2010hu . This is understood in terms of the Monte Carlo dynamics of topological modes: achieving a jump of the topological charge of the lattice gauge configuration via a local field deformation becomes increasingly more difficult as the lattice spacing gets finer due to the development of potential barriers among pseudo-topological sectors, which eventually diverge in the continuum limit to restore a proper notion of integer-valued winding number Luscher:1981zq . The rapid growth of the autocorrelation time of the topological charge thus induces a severe loss of ergodicity of the Monte Carlo Markov chain, which tends to remain frozen in a fixed topological sector for a very long time, preventing a correct sampling of the topological charge distribution. Moreover, increasing the rank of the gauge group $N$ at fixed lattice spacing further worsens the severity of freezing: in practice, at large $N$ topological freezing kicks in already at coarse lattice spacing, making it very challenging to study the continuum limit of topological quantities in this regime.
@@
While this algorithm has allowed the determination of several quantities whose calculation was either unsatisfactorily or at all attempted previously with other algorithms Hasenbusch:2017unr ; Berni:2019bch ; Bonanno:2022hmz ; Bonanno:2020hht ; Bonanno:2022yjr ; Bonanno:2023hhp ; Bonanno:2024nba ; Bonanno:2024ggk , so far only a few determinations of $\chi$ using the PTBC algorithm were provided. Some for $N=4,6$ can be found in Ref. Bonanno:2020hht , but no dedicated study of the susceptibility was performed in that paper. The goal of the present investigation is to fill this gap. I will present a dedicated high-statistics lattice study of the topological susceptibility at large $N$ using the PTBC algorithm to effectively mitigate topological freezing. Thanks to PTBC, I will simulate Yang–Mills theories for $N=4,5,6$ across the same range of lattice spacings avoiding topological freezing. The probed parameter space will include finer lattice spacings compared to previous investigations with PBC Bonati:2015sqt ; Bonati:2016tvi ; Athenodorou:2020ani ; Athenodorou:2021qvs and OBC Ce:2015qha ; Ce:2016awn , and will be similar in extension to the one used in typical $N=3$ studies. I will then combine these new large-$N$ results with the high-statistics ones I obtained for $N=3$ in Bonanno:2023ple to study the large-$N$ limit of $\chi$. Finally, I will conclude with a critical comparison of the determinations of $\chi(N)$ presented in this study with earlier ones obtained with different methods and algorithms.
@@ 2.1 Lattice action and PTBC algorithm
The calculation of $\Delta\mathcal{S}(r)$ has a negligible cost being non-vanishing only around $D$. After swaps, updating sweeps of the full lattice are alternated with some hierarchical updates concentrated around the defect $D$, which is the region where most of the new topological excitations are created/annihilated. These are chosen to have a negligible total numerical cost with respect to the full sweeps (more technical details are found in Hasenbusch:2017unr ; Bonanno:2020hht ). Each hierarchical update is followed by swap proposals too. Finally, after each series of swap proposals, a random translation of one lattice spacing in a random direction is performed on the periodic lattice, which is translation invariant. Indeed, since the region $D$ is kept fixed, performing random translations of the PBC replica effectively moves the position of the defect, so that topological excitations are created/annihilated around the lattice.
@@
The PTBC code implementation I have used for this investigation can be found in PTBC .
@@ 2.2 Determination of the topological susceptibility
$Q_{\scriptscriptstyle{\mathrm{clov}}}=\frac{1}{32\pi^{2}}\sum_{x}\sum_{\mu\nu\rho\sigma}\varepsilon_{\mu\nu\rho\sigma}\mathrm{Tr}\left[C_{\mu\nu}(x)C_{\rho\sigma}(x)\right],$ (12)
@@
It is well-known that gluonic discretizations are affected by multiplicative renormalizations Campostrini:1988cy ; Vicari:2008jw that make them non-integer on the lattice:
@@
$Q=\int\mathrm{d}^{4}x\,q(x),\qquad q(x)=\frac{1}{32\pi^{2}}\varepsilon_{\mu\nu\rho\sigma}\mathrm{Tr}\left[G_{\mu\nu}(x)G_{\rho\sigma}(x)\right],$ (16)
@@
i.e., one expects that the value of $R_{\scriptscriptstyle{\mathrm{s}}}$ should only influence the slope of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ as a function of $a^{2}$, but not its continuum limit $\chi$. In the following, I will verify this behavior by taking the continuum limit of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ at fixed value of $R_{\scriptscriptstyle{\mathrm{s}}}$ in physical units for several values of $R_{\scriptscriptstyle{\mathrm{s}}}$.
@@ 3 Numerical results
This section is devoted to the presentation of the new PTBC results for the topological susceptibility. I will start by discussing the explored parameter space and the details of the PTBC algorithm, the obtained autocorrelation times and the adopted scale setting. Then, I will present the continuum and large-$N$ extrapolations of the obtained lattice results for $\chi$. Finally, I will compare my new determinations with previous ones in the literature.
@@ 3.1 Simulation parameters, autocorrelation times and scale setting
Indeed, rescaled $\tau_{0}(Q^{2})$ can be fitted to a constant with a reduced chi-squared of $\sim 1.25$ with 10 degrees of freedom, yielding $C^{\prime}=0.081(5)$. This seems to suggest the following combined scaling towards the continuum and large-$N$ limit of the autocorrelation time of $Q^{2}$ in the periodic system:
@@ 3.2 Determination of $\chi$ at large NN
The goal of this section is to discuss the continuum and large-$N$ limits of $\chi$, and to verify the independence of $\chi$ on the choice of the smoothing radius in the continuum limit.
@@
Interestingly, lattice artifacts seem to be practically independent of $N$ for all explored values of $R_{\scriptscriptstyle{\mathrm{s}}}$—i.e., $C(R_{\scriptscriptstyle{\mathrm{s}}},N)\simeq C(R_{\scriptscriptstyle{\mathrm{s}}})$—as revealed by the plots in Fig. 5, where I compare the continuum limits of $\chi_{\scriptscriptstyle{\mathrm{L}}}$ as a function of $N$ at fixed smoothing radius. This fact was already observed in Ce:2016awn . Moreover, lattice artifacts become very small above $R_{\scriptscriptstyle{\mathrm{s}}}\sqrt{\sigma}\gtrsim 1.2$, as it can be seen from Figs. 4 and 5. For example, for the value $R_{\scriptscriptstyle{\mathrm{s}}}\sqrt{\sigma}\simeq 1.223$ reported in those figures, the coarsest lattice spacing determinations of $\chi_{{\scriptscriptstyle{\mathrm{L}}}}$ differ from their corresponding continuum ones by $\sim 2\%$ at most for all values of $N$. Since the lattice artifacts affecting $\chi_{{\scriptscriptstyle{\mathrm{L}}}}(R_{\scriptscriptstyle{\mathrm{s}}})$ have 3 different sources (the discretized lattice action used in the Monte Carlo, the discretized lattice action used in the smoothing algorithm, and the discretized observable used to compute the topological charge), the flattening of $C(R_{\scriptscriptstyle{\mathrm{s}}})$ for large enough smoothing radii could be due to an accidental cancellation among them. [^1] See Ref. Ramos:2015baa for a similar discussion about the lattice artifacts affecting the energy density computed after the gradient flow within the framework of Symanzik improvement.
@@
NN $\beta$ $a\sqrt{\sigma}$ $a^{4}\chi_{{\scriptscriptstyle{\mathrm{L}}}}\times 10^{5}$ $\chi_{{\scriptscriptstyle{\mathrm{L}}}}/\sigma^{2}$
@@ 3.4 Comparison with previous determinations of $\chi$ in the literature
0.02651(330) ETMC, Cichy et al. 15 Cichy:2015jra Wilson (Iwasaki gauge action)
@@
Table 7: Summary of continuum determinations of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(3)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 11 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action unless stated otherwise, and all assumed periodic boundaries.
@@
Table 8: Summary of continuum determinations of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(4)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 12 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 9: Summary of continuum determinations of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(5)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 13 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 10: Summary of continuum determinations of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(6)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 14 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
Table 11: Summary of continuum determinations of the topological susceptibility $\chi/\sigma^{2}$ for the $\mathrm{SU}(\infty)$ pure Yang–Mills theory. The central column translates tags appearing in Fig. 15 to the corresponding references in the bibliography. All studies adopted the standard Wilson gauge action.
@@
The case $N=3$ has the larger number of entries, and is the only case where I could retrieve determinations obtained both via gluonic definitions Durr:2006ky ; Bonati:2015sqt ; Ce:2015qha ; Athenodorou:2020ani ; Durr:2025qtq —computed employing several smoothing algorithms—and fermionic ones DelDebbio:2004ns ; Luscher:2010ik ; Cichy:2015jra ; Bonanno:2019xhg —with calculations performed adopting both chiral and non-chiral quark discretizations. Due to the much larger computational burden, fermionic definitions are systematically less precise compared to gluonic ones. Modulo this caveat, there is overall a good agreement between fermionic and gluonic results, and among all results as a whole. A naive constant fit to the full data set yields a reduced chi-squared of about 1.5 with 9 degrees of freedom, corresponding to a $\sim 15\%$ $p$-value, indeed confirming the observed agreement.